OpenStudy (anonymous):

Find the twentieth term if a36 = 78 and a42 = 90.

OpenStudy (anonymous):

From this,u get two equations......as .......a(n)=a+(n-1)d a36=a+35d=78 a42=a+41d=90

OpenStudy (anonymous):

Solve both a and d, then use a20=a+19d.find the answer!!Cheers!!

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

No probs!!